OpenStudy (anonymous):

Which equation represents the axis of symmetry of the function y = –2x² + 4x –6

OpenStudy (apoorvk):

yes man believe you can do it!!! you have a qudratic right...?? can you compare it to the std equation ax^2+bx+c=0 to find what a,b, and c are in your equation??

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{4}{2\times (-2)}=1\]